# Joint Proceedings of the ACM IUI Workshops 2024, March 18-21, 2024, Greenville, South Carolina, USA

### 26 Feb 2024 v1

@@ 2.2 Use of LLMs for Code Explanations
Generating code-explanations using LLMs requires a well-formed prompt. A verbose prompt will limit the LLM’s ability to utilize its knowledge [20]. Iterative prompts are proven to perform well [14]. A temptation to allow non-expert form input prompts will not be any good, as Zamfirescu-Pereira and colleagues [14] observed non-experts have misconceptions about LLMs and will struggle to come up with a well-formed prompt. Researchers believe that LLMs can be beneficial in environments where humans and AI can work together, where the human can perform the expert evaluation and tune the responses generated by the AI while the AI performs the time-consuming manual tasks [21].
@@ 3 The Feasibility Studies
In the third study [22], we conducted a comparative evaluation of explanations produced by experts and ChatGPT from the point of view of human users. We used two types of human users: authors (instructors and TAs) who are expected to use ChatGPT-generated explanations as the starting point in the co-creation process, and students who are the target users of the co-created product. Explanations were compared in pairs, each explanation in a pair has to be judged by completeness, and the best explanation in the pair has to be selected. A pair included an expert and a ChatGPT explanation, and the judges were not aware which source produced each explanation. The study results indicated strong preferences for ChatGPT in both groups of judges (Table 3). In general, ChatGPT explanations were rated as more complete and judged to be better in the majority of cases. However, it was not a clear win. In a substantial number of cases (15.05% for students and 27.41% for authors), expert explanations were selected as the best option in a pair.
@@
Taken the results of these two studies together, we could conclude that producing explanations for code examples is a promising application area for Human-AI co-creation. On the one hand, the LLM-generated explanations are lagging behind expert explanations in several aspects. ChatGPT explanations have higher reading difficulty than expert explanations, and they are further away from the students’ own explanations, as measured by most similarity metrics. The vocabulary data hints that ChatGPT tends to use terms, which might not be easy for the students to understand, while experts have experience in phrasing their explanations closer to the students’ active vocabulary. On the other hand, the explanations produced by ChatGPT were generally rated higher than the expert explanations by both instructors and students. These data hint that presenting ChatGPT explanations directly to students might not be a perfect solution, but they can serve as an excellent starting point for instructors in shaping their own explanations. Following that, we decided to structure the Human-AI collaboration in creating working examples as follows. Instructors have the ultimate control over producing explanations. Depending on the context (such as example complexity), they can either choose to explain example lines themselves or request AI (LLM) help in producing explanations for specific lines. In the latter case, LLM generates the initial line explanations leaving it to the instructor to accept or reject it and, if accepted, to further edit the explanation text to satisfaction. The Human-AI co-creation interface presented in the next section is based on this model of collaboration.
@@ 4 The Human-AI Co-Creation Interface Design
In the main part of the WEAT interface the problem (Figure 2-1) and the code (Figure 2-2) have to be provided by the instructor, while the explanations for each line (Figure 2-3) can be created by the instructor or generated by ChatGPT. The generated explanations could be further edited by the instructor. While we expect that co-creation of code explanations will be the preferred way to use WEAT, the system support the whole range of options from using AI explanations without human editing to creating the whole example from scratch, without the help of AI. Authors who want to start by creating explanations themselves could simply select a code line to explain (Figure 2-2) and add one or more explanation fragments to this line (Figure 2-3). The order of the fragments is important: the first fragment is displayed in PCEX when the line is clicked, while the remaining fragments can be accessed by clicking the “Additional Details” button (Figure 1-3).
@@ 5 Evaluation
After adding explanations to the example, instructor still didn’t removed the explanations for any line entirely, but removed 23 (9.7%) ChatGPT generated explanation fragments. Instructor A5 reported that he removed a number of fragments when merging two or more explanation fragments together. Since the tool did not provide support for merging fragments, it did so by copying the explanation from one fragment to the end of the other fragment and removing the obsolete fragment. In only 10 cases, instructors attempted to create new explanations from scratch, but in the end these explanations were removed. In other words, all remaining explanation fragments were originally generated by ChatGPT with some of them being edited later by the instructors. Apparently, the instructors preferred to edit the explanation fragments rather than creating them from scratch. In total, the instructors edited 66 (27.84%) of ChatGPT generated explanation fragments, on average 1.4 times (stdev=0.55). Feedback from instructors indicated that most of their edits involved summarizing, adding missing details, or removing unnecessary parts. Table 4 shows that almost half of the generated fragments were used without being touched, saving a noticeable amount of instructor time.
@@ 6 Conclusion
In this paper, we introduce a worked code example authoring tool WEAT that supports human-AI co-creation in the process of developing such examples. WEAT supports human authors by using ChatGPT for the generation of line-by-line code explanations and by providing an interface to integrate this functionality into a balanced authoring process. To the best of our knowledge, this is the first attempt to produce worked example through human-AI collaboration.
@@
An evaluation of WEAT system with five course instructors supported these expectations and provided strong evidence in favor of co-creation. As the log analysis demonstrated, in many cases, instructors choose to accept generated explanations without changes, which should have decreased the time and efforts required for example creation. Yet in other cases, the instructor rejected or edited the generated explanation to achieve the desired quality. In some cases, explanations were rejected by being simply incorrect, which stresses the importance of human presence in the authoring process. The interview with authors revealed several cases where authors acted inefficiently due to specific interface issues, such as blocking the main edit window by the generation dialog or the lack of tools to move or merge fragments. Now we are using these observations to develop an improved version of WEAT.

### 29 Feb 2024 v2

@@ 2.2 Use of LLMs for Code Explanations
Since prompts directly influence the LLM’s performance, several studies focused on exploring different prompting strategies [21, 22]. Tian et al [20] reported that a verbose prompt will limit the LLM’s ability to utilize its knowledge [20]. Iterative prompts are proven to perform well [14]. Zamfirescu-Pereira et al. [14] observed that non-experts have misconceptions about LLMs and struggle to come up with a well-formed prompt. Researchers believe that LLMs can be beneficial in environments where humans and AI can work together, where the human can perform the expert evaluation and tune the responses generated by the AI while the AI performs the time-consuming manual tasks [22].
@@ 3 The Feasibility Studies
In the third study [23], we conducted a comparative evaluation of explanations produced by experts and ChatGPT from the point of view of human users. We used two types of human users: authors (instructors and TAs) who are expected to use ChatGPT-generated explanations as the starting point in the co-creation process, and students who are the target users of the co-created product. Explanations were compared in pairs, each explanation in a pair has to be judged by completeness, and the best explanation in the pair has to be selected. A pair included an expert and a ChatGPT explanation, and the judges were not aware of which source produced each explanation. The study results indicated strong preferences for ChatGPT in both groups of judges (Table 3). In general, ChatGPT explanations were rated as more complete and judged to be better in the majority of cases. However, it was not a clear win. In a substantial number of cases (15.05% for students and 27.41% for authors), expert explanations were selected as the best option in a pair.
@@
Taking the results of these two studies together, we could conclude that producing explanations for code examples is a promising application area for Human-AI co-creation. On the one hand, the LLM-generated explanations are lagging behind expert explanations in several aspects. ChatGPT explanations have higher reading difficulty than expert explanations, and they are further away from the students’ own explanations, as measured by most similarity metrics. The vocabulary data hints that ChatGPT tends to use terms, which might not be easy for the students to understand, while experts have experience in phrasing their explanations closer to the students’ active vocabulary. On the other hand, the explanations produced by ChatGPT were generally rated higher than the expert explanations by both instructors and students. These data hint that presenting ChatGPT explanations directly to students might not be a perfect solution, but they can serve as an excellent starting point for instructors in shaping their own explanations. Following that, we decided to structure the Human-AI collaboration in creating working examples as follows. Instructors have the ultimate control over producing explanations. Depending on the context (such as example complexity), they can either choose to explain example lines themselves or request AI (LLM) help in producing explanations for specific lines. In the latter case, LLM generates the initial line explanations leaving it to the instructor to accept or reject it and, if accepted, to further edit the explanation text to satisfaction. The Human-AI co-creation interface presented in the next section is based on this model of collaboration.
@@ 4 The Human-AI Co-Creation Interface Design
In the main part of the WEAT interface, the problem (Figure 2-1) and the code (Figure 2-2) have to be provided by the instructor, while the explanations for each line (Figure 2-3) can be created by the instructor or generated by ChatGPT. The generated explanations could be further edited by the instructor. While we expect that co-creation of code explanations will be the preferred way to use WEAT, the system supports the whole range of options from using AI explanations without human editing to creating the whole example from scratch, without the help of AI. Authors who want to start by creating explanations themselves could simply select a code line to explain (Figure 2-2) and add one or more explanation fragments to this line (Figure 2-3). The order of the fragments is important: the first fragment is displayed in PCEX when the line is clicked, while the remaining fragments can be accessed by clicking the “Additional Details” button (Figure 1-3).
@@ 5 Evaluation
After adding explanations to the example, instructors still didn’t remove the explanations for any line entirely, but removed 23 (9.7%) ChatGPT generated explanation fragments. Instructor A5 reported that he removed several fragments when merging two or more explanation fragments. Since the tool did not provide support for merging fragments, it did so by copying the explanation from one fragment to the end of the other fragment and removing the obsolete fragment. In only 10 cases, instructors attempted to create new explanations from scratch, but in the end these explanations were removed. In other words, all remaining explanation fragments were originally generated by ChatGPT with some of them being edited later by the instructors. Apparently, the instructors preferred to edit the explanation fragments rather than create them from scratch. In total, the instructors edited 66 (27.84%) of ChatGPT generated explanation fragments, on average 1.4 times (stdev=0.55). Feedback from instructors indicated that most of their edits involved summarizing, adding missing details, or removing unnecessary parts. Table 4 shows that almost half of the generated fragments were used without being touched, saving a noticeable amount of instructor time.
@@ 6 Conclusion
In this paper, we introduce a worked code example authoring tool WEAT that supports human-AI co-creation in the process of developing such examples. WEAT supports human authors by using ChatGPT for the generation of line-by-line code explanations and by providing an interface to integrate this functionality into a balanced authoring process. To the best of our knowledge, this is the first attempt to develop an authoring tool that produces worked examples through human-AI collaboration.
@@
An evaluation of WEAT system with five course instructors supported these expectations and provided strong evidence in favor of co-creation. As the log analysis demonstrated, in many cases, instructors choose to accept generated explanations without changes, which should have decreased the time and effort required for example creation. Yet in other cases, the instructor rejected or edited the generated explanation to achieve the desired quality. In some cases, explanations were rejected by being simply incorrect, which stresses the importance of human presence in the authoring process. The interview with authors revealed several cases where authors acted inefficiently due to specific interface issues, such as blocking the main edit window by the generation dialog or the lack of tools to move or merge fragments. Now we are using these observations to develop an improved version of WEAT.